OpenStudy (anonymous):

Let f(x)= 1/x^2. Find formula for f`(x) using definition of derivative. Thank you!

OpenStudy (anonymous):

\[\frac{\frac{1}{(x+h)^2}-\frac{1}{(x)^2}}{h}\]

OpenStudy (anonymous):

with the lim h-->0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{\left(\frac{-h^2-2 h x}{h^2 x^2+2 h x^3+x^4}\right)}{h}=\left(\frac{-h^2-2 h x}{h^3 x^2+2 h^2 x^3+x^4h}\right)=\left(\frac{-h-2\text{ }x}{h^2 x^2+2 h x^3+x^4}\right)\]

OpenStudy (anonymous):

\[\lim_{h\to 0} \, -\frac{h+2 x}{h^2 x^2+2 h x^3+x^4}\]

OpenStudy (anonymous):

I don't understand... could you be more explicit? I'm really grateful for your help=)